OpenStudy (anonymous):

dr/dx for r=3secxcscx

zepdrix (zepdrix):

\[\Large\bf\sf r=3\sec x \csc x\]Product rule:\[\Large\bf\sf r'=\color{royalblue}{(3\sec x)'}\csc x+3 \sec x \color{royalblue}{(\csc x)'}\]

zepdrix (zepdrix):

Need to take the derivative of the blue parts. Do you remember derivative of secant and cosecant?

zepdrix (zepdrix):

If not, you can convert things to sines and cosines :o

OpenStudy (anonymous):

you have it backwards

zepdrix (zepdrix):

It's addition, there is no backwards with addition.

OpenStudy (anonymous):

product rule take derivative of the second term first and i tried changing it like that im just having trouble simplifying

zepdrix (zepdrix):

\[\Large\bf\sf r'=\color{orangered}{(3\sec x \tan x)}\csc x+3 \sec x \color{orangered}{(-\csc x \cot x)}\]Is this what you got for your derivatives?

OpenStudy (anonymous):

3secx*-cscxcotx+cscx*3secxtanx

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

where do i go from there

zepdrix (zepdrix):

Ok looks good. I guess from here you need to convert everything to sines and cosines. It should simplify down quite a bit.

OpenStudy (anonymous):

the answer is 3sec^2 x -3csc^2x so im wondering how to get there without changing to sine cosine

zepdrix (zepdrix):

You must change to sines and cosines. There might be another fancy way to do it... I'm not sure.. But this works out really nicely.

OpenStudy (anonymous):

ok well which identities are you using to get there. ?

zepdrix (zepdrix):

\[\Large\bf\sf \sec x=\frac{1}{\cos x},\qquad\qquad \csc x=\frac{1}{\sin x}\] \[\Large\bf\sf \tan x=\frac{\sin x}{\cos x},\qquad\qquad \cot x=\frac{\cos x}{\sin x}\]

OpenStudy (anonymous):

yea everything eventually cancels and you can convert back secx and cscx eh

zepdrix (zepdrix):

Well not everything, but yes there will be some nice cancellations :)

OpenStudy (kainui):

\[\frac{ 1 }{ \cos x}=(\cos x)^{-1}\] Just a nice thing to keep in mind if you hate the directionality of the quotient rule and are comfortable with the common chain rule.

OpenStudy (kc_kennylau):

Not to be confused with \(\cos^{-1}x\)